# Model-Free Assessment of Simulator Fidelity via Quantile Curves

### v1

@@ Abstract
Simulation of complex systems originated in manufacturing and queuing applications. It is now widely used for large-scale, ML-based systems in research, education, and consumer surveys. However, characterizing the discrepancy between simulators and ground truth remains challenging for increasingly complex, machine-learning–based systems. We propose a computationally tractable method to estimate the quantile function of the discrepancy between the simulated and ground-truth outcome distributions. Our approach focuses on output uncertainty and treats the simulator as a black box, imposing no modeling assumptions on its internals, and hence applies broadly across many parameter families, from Bernoulli and multinomial models to continuous, vector-valued settings. The resulting quantile curve supports confidence interval construction for unseen scenarios, risk-aware summaries of sim-to-real discrepancy (e.g., VaR/CVaR), and comparison of simulators’ performance. We demonstrate our methodology in an application assessing LLM simulation fidelity on the WorldValueBench dataset spanning four LLMs.
@@ 1 Introduction
The adoption of simulators across operations and manufacturing, agent-based modeling in social-science research, user surveys, and education/training (Zhang et al., 2019; Macal, 2016; Argyle et al., 2023; Aher et al., 2023) has accelerated with recent advances in artificial intelligence (AI). At the platform level, industry ecosystems such as NVIDIA Omniverse and Earth-2 exemplify the push toward high-fidelity, AI-enabled digital twins. Large language models (LLMs) are being applied to create and improve the ability of generative AI agents that can replicate human responses (Park et al., 2023; Lu et al., 2025). Collectively, these developments fuel the sustained growth in simulation across domains.
Against this backdrop, it becomes increasingly important to understand the discrepancy between simulation outputs and the real-world. Such discrepancies are documented across many domains: the behavior of LLMs, model discrepancy in computing systems, and sim-to-real gap in robotics (Gao et al., 2025; Durmus et al., 2023; Tobin et al., 2017; Peng et al., 2018). In light of these developments, we provide a finite-sample guaranteed distribution of the sim-to-real discrepancy without imposing any structural assumptions on the simulator.
@@
The sim-to-real discrepancy has been the focus of the uncertainty quantification (UQ) literature. Figure 1 positions our work within this landscape. Following the taxonomy of Roy and Oberkampf (2011), we distinguish input uncertainty from output uncertainty. The work on input uncertainty (see Chen et al., 2024; Barton et al., 2014; Lam, 2022) assumes that the simulator is accurate in that it creates a faithful representation of the real-world when the inputs follow the correct distribution. The goal here is to characterize how errors in the characterizing the input distribution from a finite amount of noisy input-output data propagate to errors in the output. While the initial work on input uncertainty focused on building confidence intervals for functionals of the output, recent work by Chen et al. (2024) constructs Kolmogorov-Smirnov–type uncertainty bands for the CDF of entire output discrepancy. In contrast, modern digital twins rely on complex ML models (e.g., an LLM or a deep neural network), and one does not have access to the model directly, and moreover, it is computationally prohibitive to calibrate its parameters exactly. Consequently, we treat the simulator as a black box and focus on directly characterizing the the sim-to-real discrepancy of the output, an approach termed output uncertainty quantification by Jeon et al. (2024). Moreover, the classical input uncertainty literature is primarily concerned with Monte Carlo noise, which is the variance induced by stochastic sampling conditional on a fixed input model. Our work, by contrast, explicitly incorporates both variance and bias of the simulator relative to the real world, which is precisely the focus of “sim-to-real” discrepancy. A further distinction of our work is that much of the UQ literature is asymptotic, whereas we provide finite-sample guarantees, which is critical when only a limited number of scenarios are available.
Recent work in LLM literature has also begun to focus on output uncertainty using model-agnostic discrepancy measures, e.g. Santurkar et al. (2023a) aggregate the sim-to-real gap into a single scalar summary of overall error or bias, and Huang et al. (2025) bound the discrepancy at a fixed quantile level. These approaches correspond to evaluating particular functionals of the distribution of the sim-to-real discrepancy. To obtain more flexible and informative diagnostics, we propose a model-free procedure that approximates the entire quantile function of the discrepancy. Our approach is close in spirit to conformal-inference methods (Vovk et al., 2005; Bates et al., 2021) that obtain distribution-free guarantees for a black-box predictor. However, classical conformal methods are designed for pointwise coverage at a given input rather than a characterizing the distribution of the sim-to-real discrepancy. Recent conformal variants that target distributional objects (Snell et al., 2022; Budde et al., 2025) still assume more homogeneous data structures and do not address the heterogeneous sample sizes. Our framework fills this gap by providing finite-sample, distribution-level guarantees tailored specifically to assessing black-box simulators against heterogeneous real-world data.
Contributions
Ideally, one wants to recover the quantile function for a specified discrepancy measure between the real-world and the simulated distributions. With the quantile function, we can directly compute the confidence intervals for the estimators, and risk summaries, e.g. VaR/CVaR. In practice, we only observe finite samples of the real-world and simulated outcomes, and therefore, the true quantile function is not computable. We calibrate a model-free conservative estimate for the quantile function, and prove that our estimate comes with finite-sample guarantees for any desired $\alpha$-quantile.
Building on this, we extend our framework to pairwise comparison of simulators. This allows us to benchmark a candidate simulator against a reference and to make statistically robust claims about which one is closer to reality. For instance, our procedure can certify that a given simulator achieves a smaller sim-to-real gap on at least $90\%$ of questions and pairwise hypothesis testing claims between simulators. We also illustrate our methodology on the WorldValueBench (Zhao et al., 2024) dataset, which is constructed from the World Values Survey (Haerpfer et al., 2020). We generate synthetic response profiles and compare the sim-to-real discrepancy of four different LLMs relative to the survey data.
Our main contributions are twofold:
  1. 1.
    Our procedure is model-free in that it does not impose any parametric assumptions on either the simulator or the ground truth; therefore, it can be applied to any black-box simulators. Moreover, it provides a novel assessment method to compare between simulators.
  2. 2.
    Since our procedure yields a quantile function, it naturally supports a wide range of summaries, from means to tail-risk measures such as Conditional Value at Risk (CVaR). In addition, by leveraging the simulated estimator, we can construct confidence intervals for the real-world parameters at any significance level $\alpha$.
@@ 2 Framework and Motivating Example
In this section, we start with a concrete use case that motivates our formulation, and then formally define the quantile estimation problem. Although the terminology below is specific to the example use case, the setting can be generalized. See Appendix D for more examples of simulation systems.
Suppose a media research company plans to survey customers on a particular topic with multinomial outcomes (eg, “Agree”, “Neutral”, “Disagree”). The company wants to estimate the customer opinion before committing resources for an expensive population study. The company has access to a database of past questions and human answers, and has been training an LLM-based “digital twin” for its customer base. By querying this digital twin with the new question, the company can generate an estimate of the mean response for the new question. The problem now facing the company is to characterize how close this estimate is to the true population mean. Can one provide a confidence interval for the true value? Or, better yet, estimate the quantile function for a suitable discrepancy measure between the simulator and population estimates? This work addresses precisely these uncertainty quantification problems.
To formally describe the discrepancy between simulator (LLM) and ground truth (human population), and the theoretical challenges we face, we consider two levels of randomness in this problem, which we term as the scenario uncertainty and the finite-sample uncertainty.
@@
First, scenarios (questions) are drawn as $\psi\sim\Psi$, where $\Psi$ is a scenario pool distribution, and the past dataset consists of a total of $m$ scenarios, $\{\psi_{j}\}_{j=1}^{m}\sim\Psi$. A real system (human) is characterized by a latent profile $z\in\mathcal{Z}$ with a population distribution $\mathcal{P}$ over the latent profile state $\mathcal{Z}$. For each pair $(\psi,z)$, the real system (human population) produces a categorical outcome $Y^{\mathrm{gt}},$ with conditional distribution $Q^{\mathrm{gt}}(\cdot\mid z,\psi)$ over a space $\mathcal{Y}:=\big\{y\in\mathbb{N}^{d}:\sum_{i=1}^{d}y_{i}=1\big\}$. The simulator (LLM) produces an outcome $Y^{\mathrm{sim}}$ with conditional distribution $Q^{\mathrm{sim}}(z^{\mathrm{sim}},\psi,r)$ over the same outcome space $\mathcal{Y}$, where the $z^{\mathrm{sim}}\in\mathcal{Z}_{\mathrm{sim}}$ denotes a synthetic profile drawn i.i.d. from a simulator population distribution $\mathcal{P}^{\mathrm{sim}}$ on $\mathcal{Z}_{\mathrm{sim}}$, and $r$ denotes LLM settings, including prompting strategy, hyperparameters, and other API settings. [^1] We keep $r$ fixed during calibration so that variation in $\hat{q}_{j}$ reflects scenario differences rather than encoding drift or model choice; choosing a different $r$ defines a different simulator. In the running example, $Q^{\mathrm{gt}}(\cdot|z,\psi)=\text{Categorical}(\Pi^{\mathrm{gt}}(z,\psi))$, where $\Pi^{\mathrm{gt}}(\psi,z):=\big(Q^{\mathrm{gt}}(\{1\}\mid z,\psi),\dots,Q^{\mathrm{gt}}(\{d\}\mid z,\psi)\big)\in\mathcal{P}^{d}$ and $\mathcal{P}^{d}$ is the d1d-1-simplex defined $\{u\in[0,1]^{d}:\sum_{i=1}^{d}u_{i}=1\}$. For any question $\psi$, we can marginalize the population effect, hence denote by $Q^{\mathrm{gt}}(\cdot\mid\psi)=\mathbb{E}_{z\sim\mathcal{P}}[Q^{\mathrm{gt}}(\cdot\mid z,\psi)]$, the conditional distribution of outcome $Y^{\mathrm{gt}}$ given $\psi$.
@@
Second, we only observe finite samples per scenario. For each $j\in[m]$, we are given $n_{j}$ i.i.d. profiles $z_{j,1:n_{j}}\sim\mathcal{P}$, with which we observe $n_{j}$ ground-truth outcomes $y^{\mathrm{gt}}_{j,i}\sim Q^{\mathrm{gt}}(\cdot\mid z_{j,i},\psi_{j})$. Comparably, we generate $k$ simulator outcomes $y^{\mathrm{sim}}_{j,\ell}\sim Q^{\mathrm{sim}}(\cdot\mid z^{\mathrm{sim}},\psi_{j},r)$ using a simulation pool $z^{\mathrm{sim}}_{j,1:k}\sim\mathcal{P}^{\mathrm{sim}}$ with fixed $k$ across $j$ to standardize simulator sampling. For brevity, we will write $p(\psi_{j})$ and $q(\psi_{j})$ as $p_{j}$ and $q_{j}$, and $\hat{p}_{j}$ and $\hat{q}_{j}$ be estimators of $p_{j}$ and $q_{j}$. Note that we set the number of simulated samples $k$ to be fixed across $j$ so that $\{\hat{q}_{j}\}_{j=1}^{m}$ are identically distributed. This reflects the fact that simulator sample collections are relatively inexpensive.
The dataset is $\mathcal{D}=\big\{(\psi_{j},\hat{p}_{j},\hat{q}_{j},n_{j},k)\big\}_{j=1}^{m}$. For each scenario $j$, let $\mathcal{D}^{\mathrm{gt}}_{j}=\{(z_{j,i},y^{\mathrm{gt}}_{j,i})\}_{i=1}^{n_{j}}$ denote the human-side data used to construct $\hat{p}_{j}$, and $\mathcal{D}^{\mathrm{sim}}_{j}=\{(z^{\mathrm{sim}}_{j,l},y^{\mathrm{sim}}_{j,l})\}_{l=1}^{k}$ denote the simulator outputs used to construct $\hat{q}_{j}$. The discrepancy for a specific question $\psi$ between simulated and real output distributions is defined as
@@
where $L:\Theta\times\Theta\to[0,\infty)$ is a user-chosen discrepancy function. Our method is agnostic to the choice of $L$ and allows practitioners to choose one that suits their application, such as Kullback–Leibler (KL) divergence for categorical outputs or a Wasserstein distance for nonparametric empirical distributions. Let $F_{\Delta}$ denote the cumulative distribution function (CDF) of $\Delta_{\psi}$ when $\psi\sim\Psi$, and define the associated population quantile function $V(\alpha):=\inf\{t\in\mathbb{R}:F_{\Delta}(t)\geq\alpha\},\forall\alpha\in[0,1].$
Ideally, our target is this function $V(\alpha)$, which describes the entire distribution of sim-to-real discrepancies across scenarios. However, this function is intractable, so we state our main goal:
Construct a calibrated function $\hat{V}(\cdot,\mathcal{D}):[0,1]\to\mathbb{R}$ such that, for a new $\psi\sim\Psi$ and all $\alpha\in[0,1]$,
$\mathbb{P}_{\psi\sim\Psi}\big(\Delta_{\psi}\leq\hat{V}(\alpha,\mathcal{D})\big|\mathcal{D}\big)\approx\alpha-\varepsilon_{m},$
holds with high probability over the draw of $\mathcal{D}$. The quantity $\varepsilon_{m}$ should vanish as $m\to\infty$.
Note that $\hat{V}$ is a calibrated function that is indexed both by the prescribed level $\alpha$ and by the calibration sample $\mathcal{D}$. There are two distinct sources of randomness we must control: (i) scenario uncertainty, coming from drawing finitely many scenarios $\psi_{j}\sim\Psi$, and (ii) finite-sample uncertainty within each scenario, since $p_{j}$ and $q_{j}$ are only observed through finite samples that produce $\hat{p}_{j}$ and $\hat{q}_{j}$. Our construction of $\hat{V}(\alpha,\mathcal{D})$ explicitly addresses both layers, and the main theorem will show that we obtain non-vacuous, finite-sample coverage guarantees.

### v2

@@ Abstract
As generative AI models are increasingly used to simulate real-world systems, quantifying the “sim-to-real” gap is critical. The distributional discrepancy between real and simulated outputs is a random variable driven by the stochastic input scenario. A fundamental challenge is that for any given input, the ground-truth and simulated output distributions are only observable through finite batches of samples, often of heterogeneous sizes. This renders standard predictive inference methods inapplicable, as they seek to quantify uncertainty in observable outputs rather than their underlying population parameters. To address this, we construct confidence sets for these latent parameters and use them to derive a robust proxy for the sim-to-real discrepancy. We then estimate the quantile function of this proxy to provide a comprehensive risk profile of the simulator. Our method is model-agnostic and handles general output spaces, such as categorical survey responses and continuous multi-dimensional sensor data. By rigorously accounting for sampling error, the resulting risk profile supports statistical inference for the real output distribution in a new scenario, the calculation of risk measures like Conditional Value-at-Risk (CVaR), and principled comparisons across simulators. We demonstrate the practical utility of this method by evaluating the alignment of four major LLMs with human populations on the WorldValueBench dataset.
@@ 1 Introduction
The adoption of simulators across operations and manufacturing, agent-based modeling in social-science research, user surveys, and education (Zhang et al., 2019; Macal, 2016; Argyle et al., 2023; Aher et al., 2023) has accelerated with recent advances in artificial intelligence (AI). At the platform level, industry ecosystems such as NVIDIA Omniverse and Earth-2 exemplify the push toward high-fidelity, AI-enabled digital twins, and LLMs are increasingly used to build generative agents that emulate human responses (Park et al., 2023; Lu et al., 2025). Collectively, these developments are driving growth in simulation across domains.
Against this backdrop, quantifying the distributional discrepancy between simulated and real-world outcomes is critical. This “sim-to-real” gap is a pervasive challenge in domains ranging from LLMs to computing systems and robotics (Gao et al., 2025; Durmus et al., 2023; Tobin et al., 2017; Peng et al., 2018). Crucially, this gap is not static; it is a function of the scenario (e.g., operating conditions or user prompts). Because a simulator is deployed across a heterogeneous population of scenarios, the sim-to-real gap behaves as a random variable induced by the underlying scenario distribution. Consequently, assessing fidelity requires characterizing the distribution of this random variable, rather than merely reporting a single average error. This is the primary objective of our work.
Existing uncertainty quantification methods that focus on individual outcomes (e.g., conformal prediction) are insufficient for this task, as they do not gauge how well the simulator captures the true output distribution. Conflating outcome validity with distributional fidelity can have serious downstream implications. For instance, in survey research, an LLM-based simulator might generate coherent individual responses while yielding systematically biased population-level aggregates (Westwood, 2025). Recent work has identified distributional alignment as a key objective (Cao et al., 2025), and we address the practical challenge of assessment. We provide theoretical guarantees for sim-to-real discrepancy quantification without making structural assumptions about the simulator, treating it strictly as a black box.
Contributions
We characterize the quantile function of the discrepancy between real and simulated output distributions across scenarios. Because the real-world distribution is only observed through finite, often heterogeneous, batches, the oracle quantile function is not directly computable. We construct per-scenario confidence sets for the latent real-world parameters, use them to form a proxy for the sim-to-real discrepancy, and establish finite-sample guarantees for the resulting quantile curve at any desired level. Moreover, we allow the coverage of the confidence-set to be chosen as a function of the number of sample available. This yields an adaptive procedure that produces tighter quantile curves in data-rich scenarios, while preserving reliability and interpretability when data are scarce.
We illustrate the methodology on WorldValueBench (Zhao et al., 2024), constructed from the World Values Survey (Haerpfer et al., 2020), by generating synthetic response profiles and benchmarking four LLMs against human survey data. Leveraging the large human sample size, we also study tightness and propose a data-driven procedure to construct confidence bands for the population quantile function.
Our main contributions are:
  1. (i)
    Model-free, finite-sample fidelity profiling. A black-box assessment framework with finite-sample, distributional guarantees under heterogeneous batching.
  2. (ii)
    Quantile-based risk summaries. A calibrated discrepancy quantile curve enabling the estimation of tail-risk measures, such as CVaR, and other summaries.
  3. (iii)
    Inference and comparison. Statistical inference for real-world parameters in new scenarios and principled comparisons across simulators.
@@
The sim-to-real discrepancy has been the focus of the uncertainty quantification (UQ) literature. Figure 1 positions our work within this landscape. Following the taxonomy of Roy and Oberkampf (2011), we distinguish input uncertainty from output uncertainty. Work on input uncertainty (see Chen et al., 2024; Barton et al., 2014; Lam, 2022) typically assumes that the simulator model is known and it produces a faithful representation of the real world when the input follows the correct distribution. The goal is then to characterize how errors in characterizing the input distribution, due to a finite amount of noisy data, propagate to errors in the output. While early work on input uncertainty focused on confidence intervals at a fixed significance level $\alpha$ for functionals of the output, recent work by Chen et al. (2024) constructs Kolmogorov–Smirnov–type uncertainty bands for the CDF of the output discrepancy.
In contrast, modern digital twins rely on complex ML models (e.g., an LLM or a deep neural network), and one often does not have access to the model directly. Moreover, it can be computationally prohibitive to calibrate its parameters exactly. Consequently, we treat the simulator as a black box and focus on directly characterizing the sim-to-real discrepancy of the output—an approach termed output uncertainty quantification by Jeon et al. (2024). A further distinction is that much of the UQ literature is asymptotic, whereas we provide finite-sample guarantees, which is critical when only a limited number of scenarios are available. Finally, while the classical input uncertainty literature is primarily concerned with Monte Carlo noise, which is the variance induced by stochastic sampling conditional on a fixed input model, our work incorporates both variance and bias of the simulator relative to the real world. This is precisely the focus of “sim-to-real” discrepancy.
Recent work in the LLM literature has also begun to study output uncertainty using model-agnostic discrepancy measures. For example, Santurkar et al. (2023a) aggregate the sim-to-real gap into a single scalar summary of overall error or bias, and Huang et al. (2025) bound the discrepancy at a fixed quantile level. These approaches correspond to evaluating particular functionals of the distribution of the sim-to-real discrepancy. More generally, the literature has largely emphasized scalar summaries or fixed-level bounds, leaving open the question of how to obtain a calibrated, distribution-level risk profile across scenarios.
To obtain such distribution-level guarantees in a model-agnostic way, we draw on ideas similar to conformal and predictive inference methods (Vovk et al., 2005; Bates et al., 2021), which provide finite-sample, distribution-free guarantees for black-box predictors. However, standard conformal formulations are typically stated in terms of observable outcomes (e.g., a future label or response) given each input, whereas our target is a scenario-level distributional parameter (e.g., the whole conditional output distribution or its functionals) that is only observed through finite batches, often of heterogeneous sizes. Recent conformal variants that target distributional objects (Snell et al., 2022; Budde et al., 2025) still assume homogeneous data, and do not directly address heterogeneous batching with latent scenario-level targets. We address this setting by providing finite-sample guarantees for quantile-based risk profiles of sim-to-real discrepancy when both real and simulated distributions are observed through heterogeneous batches.
@@ 2 Framework and Motivating Example
We begin with a motivating use case and then formalize the quantile estimation problem; Appendix D gives additional examples.
Suppose a media research company plans to survey customers on a particular topic with multinomial outcomes (e.g., ‘Agree’, ‘Neutral’, ‘Disagree’). The company wants to estimate customer opinion before committing resources to an expensive population study. The company is considering whether to directly use a digital twin based on pre-trained LLM or to develop its own fine-tuned LLM. In order to make this decision, the company needs to evaluate the distribution of the sim-to-real gap across the range of questions that it is likely to encounter. While pre-trained LLMs can appear to provide coherent predictions, in Appendix E we show an example where several pre-trained LLMs underperform the baseline that randomly picks answers! The methods developed in this work can help with such assessments.
@@
We formalize this setting and the associated theoretical challenges, by considering two sources of uncertainty: variation across scenarios and sampling noise within each scenario. First, we assume that scenarios (questions) are drawn from a population distribution, i.e. $\psi\sim\Psi$, and the past dataset consists of $m$ scenarios $\{\psi_{j}\}_{j=1}^{m}$ drawn i.i.d. from $\Psi$. The real system (human) is characterized by a latent profile $z\in\mathcal{Z}$ distributed according to $\mathcal{P}$ over $\mathcal{Z}$. For each pair $(\psi,z)$, the real system (human) produces a categorical outcome $Y^{\mathrm{gt}}\in\{e_{1},\cdots,e_{d}\}\subset\mathbb{R}^{d},$ with conditional distribution $Q^{\mathrm{gt}}(\cdot\mid z,\psi)$. The simulator (LLM) produces an outcome $Y^{\mathrm{sim}}$ with conditional distribution $Q^{\mathrm{sim}}(\cdot\mid z^{\mathrm{sim}},\psi,r)$ over the same outcome space, where the $z^{\mathrm{sim}}\in\mathcal{Z}_{\mathrm{sim}}$ denotes a synthetic profile drawn i.i.d. from a simulator population distribution $\mathcal{P}^{\mathrm{sim}}$ on $\mathcal{Z}_{\mathrm{sim}}$, and $r$ denotes LLM settings, including prompting strategy, hyperparameters, and other API settings. [^1] We keep $r$ fixed during calibration so that variation in $\hat{q}_{j}$ reflects scenario differences rather than encoding drift or model choice; choosing a different $r$ defines a different simulator. In the running example, $Q^{\mathrm{gt}}(\cdot|z,\psi)=\text{Categorical}(\Pi^{\mathrm{gt}}(z,\psi))$, where $\Pi^{\mathrm{gt}}(\psi,z):=\big(Q^{\mathrm{gt}}(\{1\}\mid z,\psi),\dots,Q^{\mathrm{gt}}(\{d\}\mid z,\psi)\big)\in\mathcal{P}^{d}$ and $\mathcal{P}^{d}$ is the d1d-1-simplex $\{u\in[0,1]^{d}:\sum_{i=1}^{d}u_{i}=1\}$. For any question $\psi$, we can marginalize the population effect, hence denote by $Q^{\mathrm{gt}}(\cdot\mid\psi)=\mathbb{E}_{z\sim\mathcal{P}}[Q^{\mathrm{gt}}(\cdot\mid z,\psi)]$, the conditional distribution of outcome $Y^{\mathrm{gt}}$ given $\psi$.
@@
Second, we only observe finite samples per scenario. For each $j\in[m]$, we are given $n_{j}$ i.i.d. profiles $z_{j,1:n_{j}}\sim\mathcal{P}$, with which we observe $n_{j}$ ground-truth outcomes $y^{\mathrm{gt}}_{j,i}\sim Q^{\mathrm{gt}}(\cdot\mid z_{j,i},\psi_{j})$. Comparably, we generate $k$ simulator outcomes $y^{\mathrm{sim}}_{j,\ell}\sim Q^{\mathrm{sim}}(\cdot\mid z^{\mathrm{sim}},\psi_{j},r)$ using a simulation pool $z^{\mathrm{sim}}_{j,1:k}\sim\mathcal{P}^{\mathrm{sim}}$ with fixed $k$ across $j$ to standardize simulator sampling. For brevity, we will write to be $\hat{p}_{j}$ and $\hat{q}_{j}$ be plug-in estimators of $p_{j}$ and $q_{j}$. In the motivating example these are empirical averages, though the formulation extends to more general plug-in estimators (e.g., empirical distributions or other scenario-level functionals).
Table 1: Illustrative Example of $\mathcal{D}$
Scenario True Distribution ($p_{j}$) $\boldsymbol{n_{j}}$ Real Output ($y^{\mathrm{gt}}$)
11 (0.8, 0.1, 0.1)(0.8,\ 0.1,\ 0.1) 1010 (1, 1, 2, 3, 1, 2, 1, 1, 1, 1)(1,\ 1,\ 2,\ 3,\ 1,\ 2,\ 1,\ 1,\ 1,\ 1)
22 (0.3, 0.4, 0.3)(0.3,\ 0.4,\ 0.3) 44 (3, 1, 2, 2)(3,\ 1,\ 2,\ 2)
$\vdots$ $\cdots$ $\vdots$ $\cdots$
$m$ (0.05, 0.05, 0.9)(0.05,\ 0.05,\ 0.9) 88 (1, 3, 3, 3, 3, 2, 2, 3)(1,\ 3,\ 3,\ 3,\ 3,\ 2,\ 2,\ 3)
The dataset is $\mathcal{D}=\big\{(\psi_{j},\hat{p}_{j},\hat{q}_{j},n_{j},k)\big\}_{j=1}^{m}$. For each scenario $j$, let $\mathcal{D}^{\mathrm{gt}}_{j}=\{(z_{ji},y^{\mathrm{gt}}_{ji})\}_{i=1}^{n_{j}}$ denote the human-side data used to construct $\hat{p}_{j}$, and $\mathcal{D}^{\mathrm{sim}}_{j}=\{(z^{\mathrm{sim}}_{jl},y^{\mathrm{sim}}_{jl})\}_{l=1}^{k}$ denote the simulator outputs used to construct $\hat{q}_{j}$. For concreteness, Table 1 provides an illustrative example of $\mathcal{D}$. Aligning with our motivating survey example, $p_{j}$ is a multinomial parameter and the human sample sizes $\{n_{j}\}$ may vary across scenarios, so the realized outcomes $\{y^{\mathrm{gt}}_{ji}\}$ differ both in content and in length. We record outcomes as category indices in $[d]$; equivalently, one may use one-hot encoding $e_{y}\in\mathbb{R}^{d}$. For instance, when $j=2$ the true distribution is $p_{2}=(0.3,0.4,0.3)$ and a sample of size $n_{2}=4$ yields $\hat{p}_{2}=(0.25,0.5,0.25)$.
The discrepancy for a specific question $\psi$ between simulated and real output distributions is defined as
@@
where $L:\Theta\times\Theta\to[0,\infty)$ is a user-chosen discrepancy function. We define the discrepancy in terms of $\hat{q}_{\psi}$ to reflect a fixed simulator query budget; when simulator sampling is effectively unlimited, one may replace $\hat{q}_{\psi}$ with $q_{\psi}$ to study the intrinsic discrepancy $L(p_{\psi},q_{\psi})$. We discuss this further in Section 3. Our method also works with any function $L$ and allows practitioners to choose one that suits their application, e.g. Kullback–Leibler (KL) divergence for categorical outputs or a Wasserstein distance for nonparametric empirical distributions.
Now, we can state our task:
Construct a function $\hat{V}(\cdot,\mathcal{D}):(0,1)\to\mathbb{R}$ such that, for a new $\psi\sim\Psi$ and all $\alpha\in(0,1)$,
$\mathbb{P}_{\psi\sim\Psi}\big(\Delta_{\psi}\leq\hat{V}(\alpha,\mathcal{D})\big|\mathcal{D}\big)\geq\alpha-\varepsilon_{m},$
holds with high probability over the draw of $\mathcal{D}$ and $\lim_{m\rightarrow\infty}\varepsilon_{m}=0$.
Note that $\hat{V}$ is indexed by both the prescribed level $\alpha$ and the dataset $\mathcal{D}$. Establishing our desired result requires controlling two sources of uncertainty: (i) scenario uncertainty, since we observe only finitely many scenarios $\{\psi_{j}\}_{j=1}^{m}\sim\Psi$, and (ii) finite-sample uncertainty within each scenario, since $p_{j}$ and $q_{j}$ are only observed through finite samples that produce $\hat{p}_{j}$ and $\hat{q}_{j}$. Our construction of $\hat{V}(\alpha,\mathcal{D})$ explicitly accounts for both, and the main theorem will yield non-vacuous, finite-sample coverage guarantees.